# SPLite Hand: Sparsity-Aware Lightweight 3D Hand Pose Estimation

### v2

@@ 3.1. Early Fusion
To lift 2D features into a 3D representation, we use a transformation module based on the camera’s intrinsic and extrinsic parameters. Given a 2D keypoint $(u,v)$ in camera coordinates, we associate it with a predicted depth value $d$, which allows us to compute the corresponding 3D point $(x,y,z)$ in real-world coordinates. The process consists of the following steps:
Transformation from world to camera coordinates: Using the camera extrinsic matrix, which consists of a rotation matrix $\mathbf{R}$ and a translation vector $\mathbf{T}$, we transform a point $(X_{w},Y_{w},Z_{w})$ in world coordinates to camera coordinates $(X_{c},Y_{c},Z_{c})$:
To lift 2D features into a 3D representation, we utilize a transformation module based on the camera’s intrinsic parameters. Given a 2D keypoint $(u,v)$ in camera coordinates, we associate it with a predicted depth value $d$, which allows us to compute the corresponding 3D point $(x,y,z)$ in real-world coordinates. The relationship between the 2D and 3D points is governed by the camera projection model:
$\begin{pmatrix}u\\ v\\ 1\end{pmatrix}=\mathbf{K}\begin{pmatrix}x\\ y\\ z\end{pmatrix}$
Where: - $\mathbf{K}$ is the camera’s intrinsic matrix, defined as:
$\mathbf{K}=\begin{pmatrix}f_{x}&0&c_{x}\\ 0&f_{y}&c_{y}\\ 0&0&1\end{pmatrix}$
Here, $f_{x}$ and $f_{y}$ are the focal lengths in the $x$ and $y$ directions, and $(c_{x},c_{y})$ represents the optical center (the principal point) of the camera.
The inverse of this projection operation can be used to recover the 3D coordinates $(x,y,z)$ from the 2D coordinates $(u,v)$ and the predicted depth $d$:
$\begin{pmatrix}x\\ y\\ z\end{pmatrix}=\mathbf{K}^{-1}\begin{pmatrix}u\cdot d\\ v\cdot d\\ d\end{pmatrix}$
This back-projection process assumes that we already know the depth $d$, which allows us to map the 2D coordinates back to the 3D space.
@@
Specifically, the SPLite module parallelly indexes the vertex latent, using only a quarter of the channels for the decoding convolution. This technique efficiently reduces redundant computation and memory access time while preserving model accuracy. These two features—parallel indexing and partial channel convolution—are the foundation of our SPLite decoder.

### v3

@@ 3.1. Early Fusion
To lift 2D features into a 3D representation, we utilize a feature elevation component that connects the 2D and 3D realms via a sequence of optimized procedures. This component commences with heatmap-guided estimation to forecast 2D landmark locations from the feature grid, proceeds to gather synchronized features. We associate it with a predicted depth value $d$, which allows us to compute the corresponding 3D point $(x,y,z)$ in real-world coordinates. The process consists of the following steps:
  • Heatmap-regression:
    The 2D feature grid generated by the encoder is handled to estimate 2D landmark locations. This entails producing probability distributions over a mid-resolution feature grid
    $L_{l}=\{L_{l}^{j}\}_{j=1}^{M},$
    where $M$ is the keypoint count, 21 keypoints in a hand.
  • Pose Pooling:
    Leveraging the forecasted 2D landmark locations, we derive alignment-synchronized features from the encoded 2D feature grid $F_{g}$:
    $F_{p}=\big[\,F_{g}(L_{l}^{j})\,\big]_{j=1}^{M}.$
    This produces a collection of $M$ feature vectors $F_{a}\in\mathbb{R}^{M\times C_{g}}$, where $C_{g}$ is the feature dimension, all spatially matched to the 2D landmarks.
  • Pose-to-Vertex Lifting:
    To convert the 2D-synchronized features into 3D mesh point features, we apply a trainable lifting matrix on a subsampled reference mesh. The mapping is executed as:
    $F_{\text{mesh}}=M\cdot F_{p}$
    where $M$ denotes the trained Matrix that translates landmark features to subsampled mesh features, upholding conceptual and structural integrity without depending on overt spatial transformations.
@@
Specifically, the SPLite module parallelly indexes the vertex latent, using only a quarter of the channels for the decoding convolution. This technique efficiently reduces redundant computation and memory access time while preserving model accuracy. These two features—parallel indexing and partial channel convolution are the foundation of our SPLite decoder.